# Fusion Systems related to polynomial representations of $\operatorname{SL}_{2}(q)$

### v1

@@ 5. Recognition and uniqueness results for groups and modules
The following proposition, which depends on the Classification of Finite Simple Groups, will be used to restrict the possibilities for $\mathrm{Aut}_{\mathcal{F}}(E)$ when $E$ is $\mathcal{F}$-essential.
@@
Suppose that $p$ is a prime, $X$ is a group and $T\in\operatorname{Syl}_{p}(X)$ is elementary abelian of order $p^{m}$ with $m\geq 2$. Assume that $X$ has a strongly $p$-embedded subgroup. Then $O_{p}(X)=1$ and for $K=O^{p^{\prime}}(X)=\langle T^{K}\rangle$, $\overline{K}:=K/O_{p^{\prime}}(K)$ is a non-abelian simple group and one of the following holds:
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$, $U\in\mathrm{Syl}_{p}(G)$ and $V$ is an irreducible $\mathbb{F}_{p}G$-module. Let $\mathbb{L}$ be the largest field extension of $\mathbb{F}_{p}$ over which $V$ is writable as an $\mathbb{L}$-module. Then both $C_{V}(U)$ and $V/[V,U]$ are irreducible $1$-dimensional $\mathbb{L}N_{G}(U)$-modules upon restriction.
@@
In the next result we view $V_{i}(q)$ as a module over $\mathbb{K}$. Later it will be regarded as a module over $\mathbb{F}_{p}$.
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$, $U\in\mathrm{Syl}_{p}(G)$ and $V$ is an irreducible $\mathbb{K}G$-module. If $\dim_{\mathbb{K}}V=k+1$ and $[V,U;k]\neq 1$, then $V\cong V_{k}(q)^{\theta}$ for some $\theta\in\mathrm{Aut}(\mathbb{K})$.
@@
It follows that there is a unique $10\leqslant\ell\leqslant a-$ 1 such that $d_{j}\neq 1$ and so $V=V_{d_{j}-1}(q)^{\sigma^{\ell}}$. This proves the result. ∎
@@
Suppose that $G$ has a strongly $p$-embedded subgroup, $p$ is an odd prime, $U\in\mathrm{Syl}_{p}(G)$ and $V$ is an $\mathbb{F}_{p}G$-module with $[V,O^{p}(G)]\neq 1$. If $|U|\geqslant|V/C_{V}(U)|$ or $|U|\geqslant|[V,U]|$ then $G/C_{G}(V)\cong\operatorname{SL}_{2}(|U|)$, $C_{G}(V)$ is a $p^{\prime}$-group, $V/C_{V}(G)\cong V_{1}(q)|_{\mathbb{F}_{p}}$ and $|U|=|V/C_{V}(U)|=|[V,U]|$.
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$, $q=p^{m}>p$, and $V$ is an indecomposable $\mathbb{K}G$-module with composition factors $V_{k}(q)^{\sigma^{a}}$ and $V_{\ell}(q)^{\sigma^{b}}$ where $k,\ell\in\{0,\dots,p-1\}$, $\sigma$ is the standard Frobenius automorphism of $\mathbb{K}$ defined by $x\mapsto x^{p}$ and $a\leqslant b$. Then $b=a+1$, $k=p-2$ and $\ell=1$. In particular, $\dim_{\mathbb{K}}V=p+1$.
@@
    there is $11\leqslant k,l\leqslant p-$ 1 and a submodule $W$ of $V$ such that $V/W\cong V_{k}(q)|_{\mathbb{F}_{p}}$ and $W\cong V_{l}(q)|_{\mathbb{F}_{p}}$;
@@
    if $p=3$ then for $U\in\mathrm{Syl}_{3}(G)$, $U$ does not act quadratically on $V$ and either $\dim_{\mathbb{F}_{3}}(C_{V}(U))=2m$ or $\dim_{\mathbb{F}_{3}}(V/[V,U])=2m$.
@@
It remains to consider $\mathrm{Ext}_{\operatorname{SL}_{2}(q)}^{1}(V_{k}(q)^{\sigma^{a}},V_{l}(q)^{\sigma^{b}})$. We have by Lemma 5.6 that if $\mathrm{Ext}_{\operatorname{SL}_{2}(q)}^{1}(V_{k}(q)^{\sigma^{a}},V_{l}(q)^{\sigma^{b}})$ is non-trivial then either $k=p-2,l=1$ and $b=a+1$; or $k=1$, $l=p-2$ and $b=a-1$. Furthermore, [AJL83, Corollary 4.5] implies that $\mathrm{Ext}_{\operatorname{SL}_{2}(q)}^{1}(V_{k}(q)^{\sigma^{a}},V_{l}(q)^{\sigma^{b}})$ is $1$-dimensional unless $q=9$ and $l=k=1$. If $p\neq 3$, we now have the result.
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$ where $q=p^{m}>p$ and $p$ is an odd prime, and let $S\in\mathrm{Syl}_{p}(G)$ and $11\leqslant k\leqslant p-$ 1. Then
@@
Assume that $G,V$ satisfy the hypothesis of part (1) of the lemma and that $V$ is chosen such that $V$ does not satisfy the outcome of (1) and $|V|$ is minimal. Suppose that there are $W_{1},W_{2}<C_{V}(G)$ with $|W_{1}|=p=|W_{2}|$. Consider $V/W_{i}$. Then by minimality, $C_{V/C_{V}(G)}(S)\cong C_{(V/W_{i})/C_{V/W_{i}}(G)}(S)=C_{V/W_{i}}(S)/(C_{V}(G)/W_{i})$. Hence, writing $R$ for the preimage in $V$ of $C_{V/C_{V}(G)}(S)$, we have that $[R,S]\leq W_{1}\cap W_{2}=1$ so that $C_{V}(S)/C_{V}(G)=C_{V/C_{V}(G)}(S)$, a contradiction.
@@
Let $t\in T^{\#}$ be the given element. The fact that $[V,t;k-1]=C_{V}(t),$ implies that the Jordan form of a matrix representing $t$ has $\dim_{\mathbb{L}}C_{V}(t)$ blocks each of size $k$. In particular, $t$ has order $p$, and $|V|=|C_{V}(t)|^{k}$. ∎
@@
$[V,T;k-1]=[V,r;k-1]=C_{V}(T)=C_{V}(r)$
for all $r\in T^{\#}$, then $[V,T;\ell]=[V,r;\ell]$ for all $r\in T^{\#}$ and $\ell\geq 0$. Furthermore, $[V^{*},T;k-1]=[V^{*},r;k-1]=C_{V^{*}}(T)=C_{V^{*}}(r)$ for all $r\in T^{\#}$ where $V^{*}$ is the dual of $V$.
@@
By Lemma 5.9, $|V|=|C_{V}(r)|^{k}$ and $r$ has $\dim_{\mathbb{L}}C_{V}(t)$ Jordan blocks of size $k$. The result is true for $\ell=k$ and $\ell=k-1$ by hypothesis. Suppose that it is true for $10<\ell\leqslant k-$ 1. Then $[V,T;\ell]=[V,r;\ell]$ for all $r\in T^{\#}$. By definition
$[V,T;\ell-1]=\langle[V,r;\ell-1]\mid r\in T^{\#}\rangle.$
Aiming for a contradiction, assume that $r,s\in T^{\#}$ and $[V,r;\ell-1]\neq[V,s;\ell-1]$. Then
$[V,r;\ell-1][V,s;\ell]/[V,s;\ell]\not\leq C_{V/[V,s;\ell]}(s)=[V,s;\ell-1]/[V,s;\ell]$
so that $[[V,r;\ell-1],s]\not\leq[V,s;\ell]$. Hence
$[V,s;\ell]=[V,T;\ell]\geq[[V,T;\ell-1],s][V,s;\ell]\geq[[V,r,\ell-1],s][V,s;\ell]>[V,s;\ell],$
@@
    $p$ is arbitrary, $K\cong\mathrm{PSL}_{2}(p^{m})$ and if $V$ is irreducible then $V|_{K}=V_{n}(p^{m})$ considered as a $\mathbb{F}_{p}K$-module for some even $n$ with $12\leqslant n\leqslant p-$ 1; or
@@
Suppose that Proposition 5.1(1) holds and therefore $K/Z(K)\cong\mathrm{PSL}_{2}(p^{m})$. Let $\mathbb{L}=\mathrm{End}_{\mathbb{F}_{p}}(V)$ and regard $V$ as an $\mathbb{L}G$-module so that $V$ is an absolutely irreducible $\mathbb{L}G$-module. By Lemma 5.2, $\dim_{\mathbb{L}}C_{V}(T)=1$. Since $C_{V}(t)=C_{V}(T)$ for $t\in T^{\#}$ we have $\dim_{\mathbb{L}}C_{V}(t)=1$ and consequently $V|_{\langle t\rangle}$ is indecomposable. It follows that $t$ is represented on $V$ by a matrix with a single Jordan block. Hence, writing $\dim_{\mathbb{L}}V=k+1$, we have $[V,t;k]=[V,T;k]\neq 1$ and $[V,t;k]$ is $1$-dimensional as an $\mathbb{L}$-module. Now Lemma 5.4 implies that $\mathbb{K}=\mathbb{L}$ and as a $\mathbb{L}G$-module, $V\cong V_{k}(q)^{\sigma^{\ell}}$ for some $\ell$ and $\sigma$ the standard Frobenius automorphism. Hence (1) or (2) holds in this case.
In the case of Proposition 5.1 (2), we have $K/Z(K)\cong\mathrm{Alt}(2p)$ with $p\geq 5$ and $|V|\leqslant p^{2p}$. Suppose first that $Z(K)=1$. Since $p\geq 5$, [Wag77a] together with [Jam83, Theorem 7] (see [PR02, Corollary 16.9]) imply that $V$ contains a unique non-trivial irreducible constituent isomorphic to the heart of the natural permutation module of dimension $2(p-1)$. Hence, $|V|\geqslant p^{2p-2}$ and we conclude that every $t\in T^{\#}$ has two Jordan blocks on $V$ and both are non-trivial. But then an elementary calculation shows that the $p$-cycles and the products of two $p$-cycles have fixed spaces of different dimension.
@@
by [Wag77b, 1.3 Theorem (2)]. Since $s\leqslant\log_{2}(2p)=1+\log_{2}(p)$, we have
@@
which is impossible for $p\geq 13$. For $p\in\{7,11\}$, we check that ()($\ast$)( ∗ ) does not hold. For $p=5$, $2^{.}\mathrm{Alt}(10)$ has two $8$-dimensional representations and as $|V|\leqslant 5^{10}$ and $Z(G)$ acts non-trivially on $V$, we conclude that $V$ is either irreducible or decomposable. Since $|C_{V}(s)|=5^{2}$, it is easy to see that $V$ is irreducible. We now verify by MAGMA [BCP97] that these representations do not satisfy the hypothesis of the lemma. The remaining faithful $\mathbb{F}_{5}$-representations of $2^{.}\mathrm{Alt}(10)$ have dimensions at least $28$. Thus we have no examples here.
Consider possibilities (3) and (4) so that $p=3$ and $|T|=9$. In these situations we have that $|V|\leqslant 3^{6}$ and comparing with [JLPW95] we have that $K\cong 2^{.}\mathrm{PSL}_{3}(4)$ or $K\cong\mathrm{M}_{11}$. In the former case, $K$ has a unique non-trivial module of dimension at most $6$. This module has dimension exactly $6$ and is listed as outcome (3).
In the latter case, we note that $V$ has a unique non-trivial irreducible constituent, of dimension $5$, and this constituent is either the “code” and “cocode” module for $\mathrm{M}_{11}$. We have $J\cong\mathrm{PSL}_{2}(9)$ regarded as the derived subgroup of $\mathrm{M}_{10}$ is a subgroup of $K$. We may assume that $T\leq J$ and that $(V,J)$ satisfies the hypothesis of the lemma. But then (2) holds, and we deduce that $V\cong V_{2}(9)$ which is irreducible of dimension $6$. This is impossible given that $K$ has a composition factor of dimension $5$ on $V$.
Finally consider cases (5) and (6) of Proposition 5.1. Then $\dim V\leqslant 10$. Hence $K$ must embed into $\operatorname{SL}_{10}(5)$ and, by considering the orders of the groups involved, we have $K\cong{}^{2}\mathrm{F}_{4}(2)^{\prime}$. However, by [JLPW95], the smallest characteristic $5$ representation of this group has dimension $26$. Hence these cases do not occur.
@@
We prove this by induction on $n$. For $n=1$, since $q>p$ we have that $S\leq O^{p}(H)V$ and as $C_{V}(S)<V$ we have that $[V,O^{p}(H)]\neq 1$. Observe that $|S/V|=|V/C_{V}(S)|=|C_{V}(S)|$ and so that by Lemma 5.5, we have that $C_{G}(V)=V$, $H/V\cong\operatorname{SL}_{2}(q)$ and $V$ may be viewed as a natural $\operatorname{SL}_{2}(q)$-module. Thus, the result holds when $n=1$. So assume from now that $1<n<p$.
@@
By construction, $S_{n}(q)=U\ltimes V_{n}(q)$, where $U$ is elementary abelian of order $q$ and $V_{n}(q)$ is elementary abelian, self-centralising in $S_{n}(q)$ and of order $q^{n+1}$. We aim to prove that the action of $H/V\cong\operatorname{SL}_{2}(q)$ on $V$ is irreducible, for then the result holds by Lemma 5.11. We observe that $C_{H}(V)\trianglelefteq H$ and that $[S,V]\neq 1$. Since $H/V\cong\operatorname{SL}_{2}(q)$ and $q>p$, we have that $H/C_{H}(V)\cong\operatorname{SL}_{2}(q)$ or $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$. Set $Z_{i}=Z_{i}(S)$ for $0\leq i\leq n$ where $Z_{0}(S)=1$ and $Z_{n+1}=V$.
@@
In pursuit of proving that $H/V$ is irreducible on $V$ as an $\mathbb{F}_{p}$-module, we assume aiming for a contradiction that there is $W\trianglelefteq H$ with $1\neq W<V$. Choose $W$ maximal subject to this. We either have that $W\leq C_{V}(S)=Z_{1}$, or there is some maximal $i\in\{2,\dots,n+1\}$ such that $W\cap Z_{i}\not\leq Z_{i-1}$. Then Lemma 2.10 yields $Z_{i-1}\leq W<Z_{i}$.
@@
Suppose that $i\leqslant n$ so that $W\leq Z_{n}=[V,S]$. In particular, $[V/W,S]=[V,S]/W$ has index $q$ in $V/W$. If $W=[V,S]$, then $H$ centralises $V/W$ and this is impossible as $q>p$ and $W$ was chosen maximally. Hence $W<[V,S]$. Write $\mathbb{L}$ for the largest extension of $\mathbb{F}_{p}$ for which $V/W$ is writable as an $\mathbb{L}\operatorname{SL}_{2}(q)$-module. Then by Lemma 5.2, we have that $(V/W)/[V/W,S]\cong V/[V,S]$ is irreducible as an $\mathbb{L}N_{H/V}(S/V)$-module. Hence, by Lemma 5.3 we deduce that $\mathbb{L}=\mathbb{K}$. But then $|V/W|=q^{a}$ for some $a\in\mathbb{N}$ and we conclude that $W=Z_{i}$. Now, by Lemma 2.6 we have that $S/Z_{i}\cong S_{n-i}(q)$ and by induction we have that $V/W\cong V_{n-i}(q)$. Since $|C_{V}(S)|=q$, an application of Lemma 5.8 yields that $i>1$. Hence $W=Z_{i}$ for some $i\in\{2,\dots,n-1\}$.
@@
Now, take $R<W$ with $R\trianglelefteq H$ and $R$ maximal subject to this. Then $W/R$ is an irreducible module and arguing as above, we either have that $R\leq C_{V}(S)=Z_{1}$, $Z_{i-1}\leq R$ or $R=Z_{k}$ for some $k\in\{0,1,\dots,i-2\}$ where $Z_{0}=1$. If $W\leq[V,S]$ and $Z_{i-1}\leq R$ then $W/R$ is a trivial module and, as $q>p$, we have
$Z_{i-1}<R<Z_{i}=W<Z_{n}=[V,S]$
with $V/W\cong V_{n-i}(q)$ and Lemma 5.8 implies that $C_{V/R}(S)=Z_{i+1}/R$, a contradiction to Lemma 2.10. Thus, if $W\leq[V,S]$ then $R<Z_{i-1}$.

### v2

@@ 5. Recognition and uniqueness results for groups and modules
The following proposition, which depends on the Classification of Finite Simple Groups, will be used to restrict the possibilities for $\mathrm{Aut}_{\mathcal{F}}(E)$ when $E$ is $\mathcal{F}$-essential in the fusion system $\mathcal{F}$.
@@
[…]
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$, $q=p^{m}$, $U\in\mathrm{Syl}_{p}(G)$ and $V$ is an irreducible $\mathbb{F}_{p}G$-module. Let $\mathbb{L}$ be the largest field extension of $\mathbb{F}_{p}$ over which $V$ is writable as an $\mathbb{L}$-module. Then both $C_{V}(U)$ and $V/[V,U]$ are irreducible $1$-dimensional $\mathbb{L}N_{G}(U)$-modules upon restriction.
@@
In the next result we view $V_{i}(q)$ as a module over $\mathbb{K}=\mathbb{F}_{q}$. Later it will be regarded as a module over $\mathbb{F}_{p}$.
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$, $U\in\mathrm{Syl}_{p}(G)$ and $V$ is an irreducible $\mathbb{K}G$-module. If $\dim_{\mathbb{K}}V=k+1$ and $V,U;k]\neq 1$, then $V\cong V_{k}(q)^{\theta}$ for some $\theta\in\mathrm{Aut}(\mathbb{K})$.
@@
It follows that there is a unique $0\leqslant\ell\leqslant a-1$ such that $d_{j}\neq 1$ and so $V=V_{d_{j}-1}(q)^{\sigma^{\ell}}$. This proves the result. ∎
@@
Suppose that $G$ has a strongly $p$-embedded subgroup, $p$ is an odd prime, $U\in\mathrm{Syl}_{p}(G)$ and $V$ is an $\mathbb{F}_{p}G$-module with $V,O^{p}(G)]\neq 1$. If $|U|\geqslant|V/C_{V}(U)|$ or $|U|\geqslant|[V,U]|$ then $G/C_{G}(V)\cong\operatorname{SL}_{2}(|U|)$, $C_{G}(V)$ is a $p^{\prime}$-group, $V/C_{V}(G)\cong V_{1}(q)|_{\mathbb{F}_{p}}$ and $|U|=|V/C_{V}(U)|=|[V,U]|$.
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$, $q=p^{m}>p$, and $V$ is an indecomposable $\mathbb{K}G$-module with composition factors $V_{k}(q)^{\sigma^{a}}$ and $V_{\ell}(q)^{\sigma^{b}}$ where $k,\ell\in\{0,\dots,p-1\}$, $\sigma$ is the standard Frobenius automorphism of $\mathbb{K}$ defined by $s\mapsto s^{p}$ and $a\leqslant b$. Then $b=a+1$, k=p2k=p-2 and $\ell=1$. In particular, $\dim_{\mathbb{K}}V=p+1$.
@@
    there is $1\leqslant k,l\leqslant p-1$ and a submodule $W$ of $V$ such that $V/W\cong V_{k}(q)|_{\mathbb{F}_{p}}$ and $W\cong V_{l}(q)|_{\mathbb{F}_{p}}$;
@@
    if $p=3$ then for $\in\mathrm{Syl}_{3}(G), UU$ does not act quadratically on $V$ and either $\dim_{\mathbb{F}_{3}}(C_{V}(U))=2m$ or $\dim_{\mathbb{F}_{3}}(V/[V,U])=2m$.
@@
It remains to examine when $\mathrm{Ext}_{\operatorname{SL}_{2}(q)}^{1}(V_{k}(q)^{\sigma^{a}},V_{l}(q)^{\sigma^{b}})\neq 0$. We have by Lemma 5.6 that if $\mathrm{Ext}_{\operatorname{SL}_{2}(q)}^{1}(V_{k}(q)^{\sigma^{a}},V_{l}(q)^{\sigma^{b}})$ is non-trivial then either k=p2,l=1k=p-2,l=1 and $b=a+1$; or $k=1$, l=p2l=p-2 and b=a1b=a-1. Furthermore, [AJL83, Corollary 4.5] implies that $\mathrm{Ext}_{\operatorname{SL}_{2}(q)}^{1}(V_{k}(q)^{\sigma^{a}},V_{l}(q)^{\sigma^{b}})$ is $1$-dimensional unless $q=9$ and $l=k=1$. If $p\neq 3$, we now have the result.
@@
Suppose that $G\cong\operatorname{SL}_{2}(q)$ where $q=p^{m}>p$ and $p$ is an odd prime, and let $S\in\mathrm{Syl}_{p}(G)$ and $1\leqslant k\leqslant p-1$. Then
@@
Assume that $G,V$ satisfy the hypothesis of part (1) of the lemma and that $V$ is chosen such that $V$ does not satisfy the outcome of (1) and $|V|$ is minimal. Suppose that there are $W_{1},W_{2}<C_{V}(G)$ with $|W_{1}|=p=|W_{2}|$. Consider $V/W_{i}$. Then by minimality, $C_{V/C_{V}(G)}(S)\cong C_{(V/W_{i})/C_{V/W_{i}}(G)}(S)=C_{V/W_{i}}(S)/(C_{V}(G)/W_{i})$. Hence, writing $R$ for the preimage in $V$ of $C_{V/C_{V}(G)}(S)$, we have that $R,S]\leq W_{1}\cap W_{2}=1$ so that $C_{V}(S)/C_{V}(G)=C_{V/C_{V}(G)}(S)$, a contradiction.
@@
Let $t\in T^{\#}$ be the given element. The fact that $V,t;k-1]=C_{V}(t),$ implies that the Jordan form of a matrix representing $t$ has $\dim_{\mathbb{L}}C_{V}(t)$ blocks each of size $k$. In particular, $t$ has order $p$, and $|V|=|C_{V}(t)|^{k}$. ∎
@@
$V,T;k-1]=[V,r;k-1]=C_{V}(T)=C_{V}(r)$
for all $r\in T^{\#}$, then $V,T;\ell]=[V,r;\ell]$ for all $r\in T^{\#}$ and $\ell\geq 0$. Furthermore, $V^{*},T;k-1]=[V^{*},r;k-1]=C_{V^{*}}(T)=C_{V^{*}}(r)$ for all $r\in T^{\#}$ where $V^{*}$ is the dual of $V$.
@@
By Lemma 5.9, $|V|=|C_{V}(r)|^{k}$ and $r$ has $\dim_{\mathbb{L}}C_{V}(t)$ Jordan blocks of size $k$. The result is true for $\ell=k$ and $\ell=k-1$ by hypothesis. Suppose that it is true for $0<\ell\leqslant k-1$. Then $V,T;\ell]=[V,r;\ell]$ for all $r\in T^{\#}$. By definition
$V,T;\ell-1]=\langle[V,r;\ell-1]\mid r\in T^{\#}\rangle.$
Aiming for a contradiction, assume that $r,s\in T^{\#}$ and $V,r;\ell-1]\neq[V,s;\ell-1]$. Then
$V,r;\ell-1][V,s;\ell]/[V,s;\ell]\not\leq C_{V/[V,s;\ell]}(s)=[V,s;\ell-1]/[V,s;\ell]$
so that $V,r;\ell-1],s]\not\leq[V,s;\ell]$. Hence
$V,s;\ell]=[V,T;\ell]\geq[[V,T;\ell-1],s][V,s;\ell]\geq[[V,r,\ell-1],s][V,s;\ell]>[V,s;\ell],$
@@
    $p$ is arbitrary, $K\cong\mathrm{PSL}_{2}(p^{m})$ and if $V$ is irreducible then $V|_{K}=V_{n}(p^{m})$ considered as a $\mathbb{F}_{p}K$-module for some even $n$ with $2\leqslant n\leqslant p-1$; or
@@
Suppose that Proposition 5.1(1) holds and therefore $K/Z(K)\cong\mathrm{PSL}_{2}(p^{m})$. Let $\mathbb{L}=\mathrm{End}_{\mathbb{F}_{p}}(V)$ and regard $V$ as an $\mathbb{L}G$-module so that $V$ is an absolutely irreducible $\mathbb{L}G$-module. By Lemma 5.2, $\dim_{\mathbb{L}}C_{V}(T)=1$. Since $C_{V}(t)=C_{V}(T)$ for $t\in T^{\#}$ we have $\dim_{\mathbb{L}}C_{V}(t)=1$ and consequently $V|_{\langle t\rangle}$ is indecomposable. It follows that $t$ is represented on $V$ by a matrix with a single Jordan block. Hence, writing $\dim_{\mathbb{L}}V=k+1$, we have $V,t;k]=[V,T;k]\neq 1$ and $[V,t;k]$ is $1$-dimensional as an $\mathbb{L}$-module. Now Lemma 5.4 implies that $\mathbb{K}=\mathbb{L}$ and as a $\mathbb{L}G$-module, $V\cong V_{k}(q)^{\sigma^{\ell}}$ for some $\ell$ and $\sigma$ the standard Frobenius automorphism. Hence (1) or (2) holds in this case.
In the case of Proposition 5.1 (2), we have $K/Z(K)\cong\mathrm{Alt}(2p)$ with $p\geq 5$ and $|V|\leqslant p^{2p}$. Suppose first that $Z(K)=1$. Since $p\geq 5$, [WAG77b] together with [JAM83, Theorem 7] (see [PR02, Corollary 16.9]) imply that $V$ contains a unique non-trivial irreducible constituent isomorphic to the heart of the natural permutation module of dimension 2(p1)2(p-1). Hence, $|V|\geqslant p^{2p-2}$ and we conclude that every $t\in T^{\#}$ has two Jordan blocks on $V$ and both are non-trivial. But then an elementary calculation shows that the $p$-cycles and the products of two $p$-cycles have fixed spaces of different dimension.
@@
by [WAG77a, 1.3 Theorem (2)]. Since $s\leqslant\log_{2}(2p)=1+\log_{2}(p)$, we have
@@
which is impossible for $p\geq 13$. For $p\in\{7,11\}$, we check that $(\ast)$ does not hold. For $p=5$, $2^{.}\mathrm{Alt}(10)$ has two $8$-dimensional representations and as $|V|\leqslant 5^{10}$ and $Z(G)$ acts non-trivially on $V$, we conclude that $V$ is either irreducible or decomposable. Since $|C_{V}(s)|=5^{2}$, it is easy to see that $V$ is irreducible. We now verify by MAGMA [BCP97] that these representations do not satisfy the hypothesis of the lemma. The remaining faithful $\mathbb{F}_{5}$-representations of $2^{.}\mathrm{Alt}(10)$ have dimensions at least $28$. Thus we have no examples here.
Consider possibilities (3) and (4) so that $p=3$ and $|T|=9$. In these situations we have that $|V|\leqslant 3^{6}$ and comparing with [JLP+95] we have that $K\cong 2^{.}\mathrm{PSL}_{3}(4)$ or $K\cong\mathrm{M}_{11}$. In the former case, $K$ has a unique non-trivial module of dimension at most $6$. This module has dimension exactly $6$ and is listed as outcome (3).
In the latter case, we note that $V$ has a unique non-trivial irreducible constituent, of dimension $5$, and this constituent is either the “code” and “cocode” module for $\mathrm{M}_{11}$. We have $J\cong\mathrm{PSL}_{2}(9)$, regarded as the derived subgroup of $\mathrm{M}_{10}$, is a subgroup of $K$. We may assume that $T\leq J$ and that $(V,J)$ satisfies the hypothesis of the lemma. But then (2) holds, and we deduce that $V\cong V_{2}(9)$ which is irreducible of dimension $6$. This is impossible given that $K$ has a composition factor of dimension $5$ on $V$.
Finally consider cases (5) and (6) of Proposition 5.1. Then $\dim V\leqslant 10$. Hence $K$ must embed into $\operatorname{SL}_{10}(5)$ and, by considering the orders of the groups involved, we have $K\cong{}^{2}\mathrm{F}_{4}(2)^{\prime}$. However, by [JLP+95], the smallest characteristic $5$ representation of this group has dimension $26$. Hence these cases do not occur.
@@
We prove this by induction on $n$. For $n=1$, since $q>p$ we have that $S\leq O^{p}(H)V$ and as $C_{V}(S)<V$ we have that $V,O^{p}(H)]\neq 1$. Observe that $|S/V|=|V/C_{V}(S)|=|C_{V}(S)|$ and so that by Lemma 5.5, we have that $C_{G}(V)=V$, $H/V\cong\operatorname{SL}_{2}(q)$ and $V$ may be viewed as a natural $\operatorname{SL}_{2}(q)$-module. Thus, the result holds when $n=1$. So assume from now that $1<n<p$.
@@
Recall from Notation 2.2 $S_{n}(q)=U\ltimes V_{n}(q)$, where $U$ is elementary abelian of order $q$ and $V_{n}(q)$ is elementary abelian, self-centralising in $S_{n}(q)$ and of order $q^{n+1}$. We aim to prove that the action of $H/V\cong\operatorname{SL}_{2}(q)$ on $V$ is irreducible, for then the result holds by Lemma 5.11. We observe that $C_{H}(V)\trianglelefteq H$ and that $S,V]\neq 1$. Since $H/V\cong\operatorname{SL}_{2}(q)$ and $q>p$, we have that $H/C_{H}(V)\cong\operatorname{SL}_{2}(q)$ or $H/C_{H}(V)\cong\mathrm{PSL}_{2}(q)$. Set $Z_{i}=Z_{i}(S)$ for $0\leq i\leq n$ where $Z_{0}=1$ and $Z_{n+1}=V$.
@@
In pursuit of proving that $H/V$ acts irreducibly on $V$ as an $\mathbb{F}_{p}$-module we assume, aiming for a contradiction, that there is $W\trianglelefteq H$ with $1\neq W<V$. Choose $W$ maximal subject to this. We either have that $W\leq C_{V}(S)=Z_{1}$, or there is some maximal $i\in\{2,\dots,n+1\}$ such that $W\cap Z_{i}\not\leq Z_{i-1}$. Then Lemma 2.14 yields $Z_{i-1}\leq W<Z_{i}$.
(5.14.1) $i=n+1$.
@@
Suppose otherwise so that $i\leqslant n$ and $W\leq Z_{n}=[V,S]$. In particular, $[V/W,S]=[V,S]/W$ has index $q$ in $V/W$. If $W=[V,S]$, then $H$ centralises $V/W$ and this is impossible as $q>p$ and $W$ was chosen maximally. Hence $W<[V,S]$. Write $\mathbb{L}$ for the largest extension of $\mathbb{F}_{p}$ for which $V/W$ is writable as an $\mathbb{L}\operatorname{SL}_{2}(q)$-module. Then by Lemma 5.2, we have that $(V/W)/[V/W,S]\cong V/[V,S]$ is irreducible as an $\mathbb{L}N_{H/V}(S/V)$-module. Hence, by Lemma 5.3 we deduce that $\mathbb{L}=\mathbb{K}:=\mathbb{F}_{q}$. But then $|V/W|=q^{a}$ for some $a\in\mathbb{N}$ and we conclude that $W=Z_{i}$. Now, by Lemma 2.10 we have that $S/Z_{i}\cong S_{n-i}(q)$ and by induction we have that $V/W\cong V_{n-i}(q)$. Since $|C_{V}(S)|=q$, an application of Lemma 5.8 yields that $i>1$. Hence $W=Z_{i}$ for some $i\in\{2,\dots,n-1\}$.
@@
Now, take $W_{0}<W$ with $W_{0}\trianglelefteq H$ and $W_{0}$ maximal subject to this. Then $W/W_{0}$ is an irreducible module and arguing as above, we either have that $W_{0}\leq C_{V}(S)=Z_{1}$, $Z_{i-1}\leq W_{0}$ or $W_{0}=Z_{k}$ for some $k\in\{0,1,\dots,i-2\}$ where $Z_{0}=1$. If $Z_{i-1}\leq W_{0}$ then $W/W_{0}$ is a trivial module and, as $q>p$ and $W_{0}$ is maximal, we have
$Z_{i-1}<W_{0}<Z_{i}=W<Z_{n}=[V,S]$
with $V/W\cong V_{n-i}(q)$ and Lemma 5.8 implies that $C_{V/W_{0}}(S)=Z_{i+1}/W_{0}$, a contradiction to Lemma 2.14. Thus, $W_{0}<Z_{i-1}$.